Find the value of tan



-11 + tan-12 + tan-13.

Grade:12

1 Answers

Moni RS
14 years ago
let tan-1(1) + tan-1(2) + tan-1(3) = x but tan-1(1)= p/4 tan-1(2)+tan-1(3)=tan-1((2+3)/(1-2*3))=tan-1(-1)= - tan-1(1) therefore tan-1(1)+(- tan-1(1))=0.